# existence1.bib

### v1

@@ 6.2 Constraint Set given by Functional Inequalities
In this section, we consider problems of the following form:
@@
where $C$ is a closed set, and $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$and each $g_{j}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ are proper closed functions. Existence of solutions to this problem has been studied in both general settings [ozdaglar_tseng_2006] as well as special settings [Belousov_Klatte_2002, Eaves-quad-programming, Frank_Wolfe_Theorem_56, Luo_Zhang_1999, Terlaky-LP].
We provide conditions for the existence of solutions based on Theorem 4.2.2 and Theorem 4.2.3 combined.
Theorem 6.2.1.
Let $C$ be a closed set, and let $f$ and each $g_{j}$ be proper closed functions with $C\cap\left(\cap_{j=1}^{m}{\rm dom}g_{j}\right)\cap{\rm dom}f\neq\emptyset$. Let $X=\{x\in C\mid g_{j}(x)\leq 0,j=1,\ldots,m\}$, and assume that $f$ exhibits asymptotically bounded decay on the set $X$ with respect to $g(x)=\|x\|^{p}$ for some $p\geq 0$ (or with respect to a coercive function $g$ satisfying the assumptions of Theorem 4.2.3). Then, the problem (23) has a finite optimal value and an optimal solution exists under any of the following conditions:
(C1)  $C_{\infty}\cap(\cap_{j=1}^{m}(L_{0}(g_{j}))_{\infty})\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f),$
(C2)  $C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j}))\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f).$
@@
By Proposition 2.2.1, we have that
@@
Thus, if the condition (C1) holds, then it follows that $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ and the result follows by Theorem 4.2.2 (or Theorem 4.2.3). If condition (C2) holds, then by Proposition 2.3.1 we have
$(L_{0}(g_{j}))_{\infty}\subseteq\mathcal{K}(g_{j})\qquad\hbox{for all }j=1,\ldots,m.$
By combining these relations with (25), again we have that $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ and the result follows as in the preceding case. ∎∎
In Theorem 6.2.1, we could not write the cone $\mathcal{R}(X)$ in terms of such cones of the individual sets defining the set $X$, as there is no particular rule that can be applied here, in general. In the case when $X$ and the functions $g_{j}$ are convex, the conditions (C1) and (C2) of Theorem 6.2.1 coincide, since $(L_{0}(g_{j}))_{\infty}=\mathcal{K}(g_{j})$ for all $j$ by Proposition 2.3.1.
@@
Theorem 6.2.2.
Let assumptions of Theorem 6.2.1 hold. Additionally, assume that the set $C$ is convex and that each $g_{j}$ is a convex function. Then, the problem (23) has a finite optimal value and an optimal solution exists under any of the following conditions:
(C3)  $C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j}))\cap\mathcal{K}(f)\subseteq\mathcal{R}(C)\cap(\cap_{j=1}^{m}\mathcal{R}(L_{0}(g_{j}))\cap\mathcal{R}(f),$
(C4)  $C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j}))\cap\mathcal{K}(f)\subseteq{\rm Lin}(C)\cap(\cap_{j=1}^{m}\mathcal{C}(g_{j}))\cap\mathcal{R}(f).$
@@
Let the condition (C3) hold. Since $X$ is convex and each $g_{j}$ is convex, we have that $X_{\infty}=C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j})).$ Moreover, by Proposition 3.2.2 we have
@@
Thus, the condition (C2) of Theorem 6.2.1 is satisfied and the result follows. Suppose that the condition (C4) holds. Then, since
@@
it follows that the condition (C2) of Theorem 6.2.1 is satisfied. ∎∎
Now consider problem (23) where each $g_{j}$ is a convex polynomial. As a corollary of Theorem 6.2.2, we have the following result which extends Theorem 3 in [Belousov_Klatte_2002], where $f$ is also convex polynomial.
Corollary 6.2.1.
Consider the problem (23), where $C=\mathbb{R}^{n}$, the objective function $f$ is proper and closed, and each $g_{j}$ is a convex polynomial. Further, assume that
@@
Then, the problem has a finite optimal value and a solution exists.
@@
By Example 3.4, a polynomial of order $p$ asymptotically decays with respect to the function $g(x)=\|x\|^{p}$. Since each $g_{j}$ is a polynomial, by Lemma 3.3.2 we have that $\mathcal{R}(g_{j})=\mathcal{C}(g_{j})$ for all $j$. Therefore, the condition (C4) of Theorem 6.2.2 is satisfied, with $C=\mathbb{R}^{n}$, and the result follows. ∎∎
The following example shows that when the condition on the asymptotic cones of the functions $g_{j}$ and $f$ of Corollary 6.2.1 is violated, the problem (23) may not have a solution.
Example 6.2 (Example 2 in [Luo_Zhang_1999]).
@@
The objective polynomial $f$ is non-convex while both $g_{1}$ and $g_{2}$ are convex polynomials. By the convexity of the constraint sets, we have
@@
Moreover, by Lemma 3.3.2 we have $\mathcal{C}(g_{j})=\{0\}$ for $j=1,2$ so that $\mathcal{C}(g_{1})\cap\mathcal{C}(g_{2})=\{0\}.$ The function $f$ is a polynomial of order 2, so by Lemma 3.3.1 we have $\mathcal{K}(f)=\{d\mid f(d)\leq 0\}.$ Then, it follows that
@@
does not hold, since a nonzero direction $d=(0,0,d_{3},0)$ with $d_{3}>0$ belongs to $\mathcal{K}(g_{1})\cap\mathcal{K}(g_{2})\cap\mathcal{K}(f)$ but not to $\mathcal{C}(g_{1})\cap\mathcal{C}(g_{2})\cap\mathcal{R}(f)$. However, the optimal value of the problem is $f^{*}=-1$ which is not attained, as shown in [Luo_Zhang_1999]. ∎
We conclude this section by comparing our results with the main result of [ozdaglar_tseng_2006], which is Proposition 3.1 therein, and we demonstrate on an example that our results are more general.
Consider the following problem
minimize $f(x_{1},x_{2})=x_{1}^{2}+x_{2}$ (27)
subject to $g_{1}(x_{1},x_{2})=\sqrt{|x_{1}|}-x_{2}\leq 0.$ (28)
The optimal value of the problem is $f^{*}=0$ and the optimal point is (0,0). The objective function is a convex polynomial, which exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{2}$. By Lemma 3.3.1 we have $\mathcal{K}(f)=\{(0,d_{2})\mid d_{2}\leq 0\}$, and $\mathcal{R}(f)=\{(0,0)\}$ by Lemma 3.3.2. For the set $X=\{x\in\mathbb{R}^{2}\mid\sqrt{|x_{1}|}\leq x_{2}\}$, we have $X_{\infty}=\{(d_{1},d_{2})\mid d_{1}\in\mathbb{R},d_{2}\geq 0\}$ and $\mathcal{R}(X)=X_{\infty}$. Hence, $X_{\infty}\cap\mathcal{K}(f)=\{(0,0)\}$ and $\mathcal{R}(X)\cap\mathcal{R}(f)=\{(0,0)\}$, implying that the conditions of Theorem 4.2.2 are satisfied. Hence, by Theorem 4.2.2, the problem has a finite optimal value and a solution exists.
Proposition 3.1 of [ozdaglar_tseng_2006] assumes that the optimal value of the problem (27) is 0, which is the case. Proposition 3.1 requires that $\{x\mid f(x)\leq 0\}$ is contained in the domain of the function $g_{1}(x_{1},x_{2})=\sqrt{|x_{1}|}-x_{2}$ (assumption (A1)(b) of [ozdaglar_tseng_2006]) and that $L_{\gamma}(f)\cap L_{0}(g_{1})\neq\emptyset$ for all $\gamma>0$, which are both satisfied.
Additionally, Assumption (A1)(a) of [ozdaglar_tseng_2006] needs to be satisfied, which requires some special notions. One of them is an asymptotically nonpositive direction $d$ of a function $h$ requiring that for a sequence $\{x_{k}\}\subset{\rm dom}(h)$, with $\|x_{k}\|\to\infty$ and $x_{k}/\|x_{k}\|\to d$, it holds that
$\limsup_{k\to\infty}h(x_{k})\leq 0.$
Another one is a direction $d$ along which the function $h$ recedes below 0 on a set $S\subset{\rm dom}h$ requiring that
for $\hbox{for every $x\in S$ there is $\bar{\alpha}\geq 0$ such that }\ f(x+\alpha d)\leq 0\quad\hbox{ for all }\alpha\geq\bar{\alpha}.$
Lastly, for an asymptotically nonpositive direction $d$ of a function $h$, it is said that $h$ retracts along $d$ on a set $S\subset{\rm dom}h$ if for any sequence $\{x_{k}\}\subset S$, with $\|x_{k}\|\to\infty$ and $x_{k}/\|x_{k}\|\to d$, there exists $\bar{k}$ such that
$f(x_{k}-d)\leq\max\{0,f(x_{k})\}\qquad\hbox{for all }k\geq\bar{k}.$
Assumption (A1)(a) as applied to problem (27) with a single constraint, requires that the following conditions hold:
  • (D1)
    For every asymptotically nonpositive direction $d$ of $f$, the function $f$ either (i) recedes below 0 along $d$ on ${\rm dom}f$ or (ii) retracts along $d$ on ${\rm dom}f$.
  • (D2)
    For every asymptotically nonpositive direction $d$ of $g_{1}$, the function $g_{1}$ either (i) recedes below 0 along $d$ on its domain or (ii) retracts along $d$ on its domain and recedes below 0 on the level set $L_{0}(g_{1})$.
If a direction $d$ is asymptotically nonpositive direction of a function $h$, then we must have $h_{\infty}(d)\leq 0$, which follows from the definition of the asymptotic function. Since $\mathcal{K}(f)=\{(0,d_{2})\mid d_{2}\leq 0\}$, we can see that any $d\in\mathcal{K}(f)$ is asymptotically nonpositive direction. Thus, $f$ satisfies the condition (D1).
Next, consider the constraint function $g_{1}$ and a sequence $x_{k}=(k,\sqrt{k})$ for $k\geq 1.$ We have $\|x_{k}\|\to\infty$, $x_{k}/\|x_{k}\|\to(1,0)$, and $g_{1}(x_{k})=0$ for all $k$. Thus, $d=(1,0)$ is an asymptotically nonpositive direction of $g_{1}$. However, the function does not recede below 0 neither on ${\rm dom}g_{1}$ nor on $L_{0}(g_{1})$. To see this, note that $x=0$ belongs to $L_{0}(g_{1})$ and ${\rm dom}g_{1}$, while for any $\alpha>0$,
$g_{1}(0+\alpha d)=g_{1}(\alpha,0)=\sqrt{\alpha}>0.$
Thus, $g_{1}$ does not satisfy the condition (D2) so Proposition 3.1 of [ozdaglar_tseng_2006] cannot be used to assert the existence of solutions to problem (27), while our Theorem 4.2.2 can be used.

### v3

@@ 6.2 Constraint Set given by Functional Inequalities
In this section, we consider a problem of the following form:
@@
where $C$ is a closed set, and $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$and each $g_{j}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is a proper closed function. Existence of solutions to this problem has been studied in both general settings [16] as well as in special settings [5, 9, 11, 14, 20]. To the best of our knowledge all the results provided in this section are new, and more general than those in [5, 9, 11, 14, 20, 16], which only provide sufficient conditions.
We provide necessary and sufficient conditions for the existence of solutions to problem (19) based on Theorems 4.24.5.
Theorem 6.2 (Nonconvex function and nonconvex constraint set).
Let $C$ be a closed set, and let $f$ and each $g_{j}$ be proper closed functions with $C\cap\left(\cap_{j=1}^{m}{\rm dom}g_{j}\right)\cap{\rm dom}f\neq\emptyset$. Let $X=\{x\in C\mid g_{j}(x)\leq 0,j=1,\ldots,m\}$, and assume that $f$ exhibits asymptotically bounded decay on the set $X$ with respect to a coercive function $g$ satisfying the assumptions of Theorem 4.2, or with respect to $g(x)=\|x\|^{p}$ for some $p\geq 0$. Then, either one of the following two conditions
(C1)  $C_{\infty}\cap(\cap_{j=1}^{m}(L_{0}(g_{j}))_{\infty})\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$
(C2)  $C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j}))\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$
is necessary and sufficient for the problem (19) to have a finite optimal value and an optimal solution.
@@
By Proposition 2.1, we have that
@@
Thus, if condition (C1) holds, then $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ and problem (19) has a finite optimal value and an optimal solution by Theorem 4.2 (or Theorem 4.3).
If condition (C2) holds, then by Proposition 2.3 we have $(L_{0}(g_{j}))_{\infty}\subseteq\mathcal{K}(g_{j})$ for all $j=1,\ldots,m.$ By combining these relations with (21), again we have that $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ and the result follows as in the preceding case.
The necessity of these conditions follows by Theorem 4.5. ∎
In Theorem 6.2, we could not write the cone $\mathcal{R}(X)$ in terms of such cones of the individual sets defining the set $X$, as there is no particular rule that can be applied here, in general. When $X$ and the functions $g_{j}$ are convex, the conditions (C1) and (C2) of Theorem 6.2 coincide, since $(L_{0}(g_{j}))_{\infty}=\mathcal{K}(g_{j})$ for all $j$ by Proposition 2.3.
@@
Theorem 6.3 (Nonconvex function with convex constraint set).
Let assumptions of Theorem 6.2 hold. Additionally, assume that the set $C$ is convex and that each $g_{j}$ is a convex function. Then, either one of the following two conditions
(C3)  $C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j}))\cap\mathcal{K}(f)\subseteq\mathcal{R}(C)\cap(\cap_{j=1}^{m}\mathcal{R}(L_{0}(g_{j}))\cap\mathcal{R}(f)$
(C4)  $C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j}))\cap\mathcal{K}(f)\subseteq{\rm Lin}(C)\cap(\cap_{j=1}^{m}\mathcal{C}(g_{j}))\cap\mathcal{R}(f)$
is necessary and sufficient for problem (19) to have a finite optimal value and an optimal solution.
@@
Let the condition (C3) hold. Since $X$ is convex and each $g_{j}$ is convex, we have that $X_{\infty}=C_{\infty}\cap(\cap_{j=1}^{m}\mathcal{K}(g_{j})).$ Moreover, by Proposition 3.2 we have
@@
Thus, the condition (C2) of Theorem 6.2 is satisfied and the result follows. Suppose that the condition (C4) holds. Then, since
@@
the condition (C2) of Theorem 6.2 is satisfied. ∎
Now consider problem (19) where each $g_{j}$ is a convex polynomial. As a corollary of Theorem 6.3, we have the following result.
Corollary 6.1 (Nonconvex polynomial function with convex polynomial constraints).
Consider the problem (19), where $C=\mathbb{R}^{n}$, the objective function $f$ is proper and closed, and each $g_{j}$ is a convex polynomial. Then, the problem has a finite optimal value and a solution exists if and only if
@@
By Example 3.4, a polynomial of order $p$ asymptotically decays with respect to the function $g(x)=\|x\|^{p}$. Since each $g_{j}$ is a polynomial, by Lemma 3.2 we have that $\mathcal{R}(g_{j})=\mathcal{C}(g_{j})$ for all $j$. Therefore, the condition (C4) of Theorem 6.3 is satisfied, with $C=\mathbb{R}^{n}$, and the result follows. ∎
Corollary 6.1 extends Theorem 3 of [5] in two ways. Firstly Theorem 3 applies to the convex case as $f$ is a convex polynomial. Second Theorem 3 provides only sufficient conditions.
The following example shows that when the condition of Corollary 6.1 is violated, problem (19) may not have a solution.
Example 6.2 (Example 2 in [14]).
@@
The objective polynomial $f$ is nonconvex while both $g_{1}$ and $g_{2}$ are convex polynomials.
By the convexity of the constraint sets, we have
@@
Moreover, by Lemma 3.2 we have $\mathcal{C}(g_{j})=\{0\}$ for $j=1,2$ so that $\mathcal{C}(g_{1})\cap\mathcal{C}(g_{2})=\{0\}.$ The function $f$ is a polynomial of order 2, so by Lemma 3.1 we have $\mathcal{K}(f)=\{d\mid f(d)\leq 0\}.$ Then, it follows that
@@
does not hold, since a nonzero direction $d=(0,0,d_{3},0)$ with $d_{3}>0$ belongs to $\mathcal{K}(g_{1})\cap\mathcal{K}(g_{2})\cap\mathcal{K}(f)$ but not to $\mathcal{C}(g_{1})\cap\mathcal{C}(g_{2})\cap\mathcal{R}(f)$. However, the optimal value of the problem is $f^{*}=-1$ which is not attained, as shown in [14]. ∎
@@
Acknowledgements This work has been supported by the ONR award N00014-21-1-2242 and the NSF award CIF 2134256.